OpenStudy (anonymous):

Determine the solution set of (x - 3)^2 = 49.

OpenStudy (anonymous):

(x - 3)^2 = 49 (x - 3) = 7 or (x - 3) = -7

OpenStudy (anonymous):

a. {6, -8} b. {4, 10} c. {10, -4}

OpenStudy (anonymous):

(x - 3) = 7 or (x - 3) = -7 x=10 or x= -4